OpenStudy (anonymous):

Could somebody show me how to simplify this?(5√2-1)²

OpenStudy (mayankdevnani):

welcome to openstudy!

OpenStudy (anonymous):

Thanks! :)

OpenStudy (mayankdevnani):

did you try it ?

OpenStudy (anonymous):

Yeah, I tried to simplify it myself, but I'm not entirely sure where to start.

OpenStudy (anonymous):

Do I get the √2 first before I try and solve it?

OpenStudy (mayankdevnani):

use \[\large \bf (a-b)^2=a^2+b^2-2ab\]

OpenStudy (anonymous):

Ahh, so I'm missing a formula, makes sense, lol. Thanks! I can take it from here.

OpenStudy (mayankdevnani):

good !

OpenStudy (anonymous):

Well, I thought I could figure it out with the formula... Sorry, I have not a single clue what to do with that formula. There is no a or b in (5√2-1)² could you show me how to apply the formula to the question please?

OpenStudy (mayankdevnani):

\[\large \bf (5 \sqrt2)^2+(1)^2-2(5\sqrt{2})(1)\]

OpenStudy (phi):

you don't need a formula, but it helps to memorize them if you want to be "fast" the little 2 in the upper right means "square" or multiply by itself \[ ( 5 \sqrt{2} -1 ) ( 5 \sqrt{2} -1 ) \] have you learned FOIL to multiply "binomials"?

OpenStudy (phi):

the F in FOIL means "first" which is to remind you to multiply the two "first" terms: \[ 5 \cdot \sqrt{2} \cdot 5 \cdot \sqrt{2} \\= 5 \cdot 5 \cdot \sqrt{2} \cdot \sqrt{2} \]